If a parallel beam of light of energy density U falls normally on an object and totally absorbed what is the pressure exerted on body
1
Expert's answer
2019-06-20T11:16:47-0400

photon momentum equals to

"p_p = \frac{\epsilon}{c},"

therefore by unit area per unit of time,

"\frac{U}{c}"

momentum is absorbed; this value equals to pressure p, that is

"p =\frac{U}{c}"